### v1

# Diffusion-Reinforcement Learning Hierarchical Motion Planning in Adversarial Multi-agent Games

Abstract
Reinforcement Learning- (RL-)based motion planning has recently shown the potential to outperform traditional approaches from autonomous navigation to robot manipulation. In this work, we focus on a motion planning task for an evasive target in a partially observable multi-agent adversarial pursuit-evasion games (PEG). These pursuit-evasion problems are relevant to various applications, such as search and rescue operations and surveillance robots, where robots must effectively plan their actions to gather intelligence or accomplish mission tasks while avoiding detection or capture themselves. We propose a hierarchical architecture that integrates a high-level diffusion model to plan global paths responsive to environment data while a low-level RL algorithm reasons about evasive versus global path-following behavior. Our approach outperforms baselines by 51.2% by leveraging the diffusion model to guide the RL algorithm for more efficient exploration and improves the explanability and predictability. [^1] Code and implementation details are at https://github.com/CORE-Robotics-Lab/Opponent-Modeling
I Introduction
Multi-agent pursuit-evasion games (PEG) [1] are ubiquitous in real-world robotics scenarios where robotic systems could be applied to tracking drug smugglers [2], defensive escorting [3], or robot soccer [4]. In PEG, one team seeks to track its target while the target agent or team employs evasive maneuvers, and research naturally focuses on developing methods for either the pursuers [5] or evaders [6, 7, 8]. While there has been extensive work on developing search policies, including robust tracking frameworks [9], topology map maintenance [10], and multi-hypothesis filters [11], relatively little attention has been paid to evasion. This lack of emphasis provides an opportunity to design a novel evasive policy which is also useful in the real-world scenarios (e.g. business ship evading from pirates).
Traditional approaches to develop evasion policies typically assume full environment observability and are designed to optimize capture time [12] or an ad hoc utility function [13].
Figure 1: Diffusion-RL framework overview: We first collect RRT* paths into a dataset. Then we use a diffusion model to learn the distribution of the RRT* path and generate samples as the high-level global plans to help learn a low-level evasive RL policy. A posterior costmap is built based on the learned hierarchy and detection risk and can be used to select the best global path in the inference stage.
More recently, learning-based methods, such as Graph Neural Networks [14], diffusion models [15] or reinforcement learning (RL) methods [16] have been widely used in the pursuit-evasion games and in similar navigation problems [17, 18, 19] (e.g. goal-reaching, path-planning, etc.). These methods often outperform expert policies [20]. Unfortunately, learning the RL policy is usually subject to two key challenges—partial observability and high sample complexity when exploring in large-scale, partially observable environments, as considered in our work.
Researchers propose hierarchical approaches to help exploration by structuring the goal, policies, or rewards. Goal-oriented reinforcement learning (e.g. hindsight experience replay [21]) increases the reward density by assuming the goal has been reached in each exploration trajectory and correcting observation and reward accordingly. Hierarchical reinforcement learning [22] improves the sample efficiency from spatial or temporal abstraction by multiple hierarchically organized RL policies, and hybrid reward structure based methods [23] try to capture the multi-modal distribution of reward where each part of value function can be a low-dimensional representation. However, their applications are either tested in fully observable games [22] or single agent scenarios [24], and sometimes require expert prior knowledge as primitives [25].
In this paper, we propose a novel hierarchical learning framework that utilizes diffusion and RL as high and low level policies respectively. The diffusion model is used as a global path generator to satisfy static map constraints, such as terminal states and obstacle avoidance. Additionally, we employ a low-level RL policy that learns to follow the waypoints while prudently evading captures (see Figure 1). Our key insight is that using the diffusion model can improve the RL performance by constraining its exploration to high-value state-action regions. We find our method is especially useful for the evader to learn escaping behaviors in the large multi-agent, multi-goal partially observable settings — Prisoner Escape and Narco Interdiction Domain (see §III-B). Increasing understanding of potential evader strategies will help future development of countermeasures.
Contributions: In summary, our contributions are three-fold:
  1. 1.
    Our work is the first to create a learning-based framework for evaders in large partially observable environments that learns without expert domain knowledge.
  2. 2.
    We propose a novel hierarchical system consisting of a diffusion model as a high level global path planner and a low level RL agent to learn evasive maneuvers, which significantly outperforms all baselines.
  3. 3.
    To the best of our knowledge, we pioneer the use of task-oriented costmap construction to enhance explainability, predictability, and flexibility, thereby facilitating a deeper understanding of the evader’s performance.
II Related Work
II-A Pursuit-Evasion Games
A perfect information PEG (i.e., full observability) can be modeled as a differential game and solved using control theory [26, 27, 28, 29]. In comparison, the incomplete information PEG (i.e., partial observability) is a more realistic setting. Researchers commonly use probabilistic models [30, 31] and formulate PEG as optimization problems with the objective of maximizing the payoff [32, 33]. Therefore, recent research also focuses on deriving the optimal policy via RL [34, 32, 35]. Unfortunately, most of RL work focuses on pursuit policies [34, 32, 36]. While Generative Adversarial Imitation Learning has been employed for training an evader policy, this approach requires a dataset showing expert evasive behaviors and a complex curriculum RL scheme [35]. Furthermore, prior work only validates this approach on a small map. In comparison, our method trains evasive RL policy from scratch without any evasion behavior demonstrations and is simpler and more suitable for large, adversarial, and partially observable domains.
II-B Diffusion Models in Path Generating
Diffusion models are a class of generative models that have demonstrated impressive performance in generating sequential data across several applications, such as image synthesis and video generation [37, 38, 39]. Recently, there have been several works employing diffusion models in trajectory generation tasks. Diffuser [40] trains a diffusion model for trajectory generation on offline datasets and plans future trajectories with guided sampling. Diffusion policy extended this work to imitation learning and was able to learn robust policies for various pushing tasks conditioned only on visual inputs [41]. Recent work [15] has extended these approaches for generating trajectories of multiple opponents. There have also been many following works using diffusion models to synthesize trajectories and augment experience buffer for efficient offline RL training [42, 43]. However, these works only function in fully-observable domains, where building a world model is feasible. We instead use the diffusion model as a trajectory planner that allows the reinforcement learning agent to generate diverse plan options in our partially observable pursuit-evasion domain.
II-C Reinforcement Learning in Motion Planning
Recent works have demonstrated the success of RL for motion planning problems in dynamic environments [44, 45, 46, 47, 17]. Many papers discuss how to use RL to avoid obstacles by analyzing exteroceptive sensor information such as RGB cameras [44, 47], LIDARs [46, 47], or laser data converted from images [45]. However, none of these prior works assumes the an adversarial environment, where the “obstacles” are intelligent pursuers, and the map size is as large as in our scenario.
The most similar work from the system design perspective is from [17] which also employs a hierarchical structure in goal-conditioned planning. However, their environment does not incorporate adversaries or multiple goals. Additionally, they require expert datasets for training their RL policy. Dynamic obstacles are considered in [48]; however, these obstacles can only run on a pre-defined segment with a constant velocity, and all the environment objects are observable such that A* search can be employed effectively. In our scenario, we assume the cameras are unobservable to the evader, while the pursuit team can leverage these cameras to detect and track the evader. Therefore, we use a diffusion model to generate diverse global candidate paths and make it possible for RL to do useful exploration about the map.
III Environment
Figure 2: Prisoner Escape Domain (left) and Narco Traffic Interdiction Domain (right)
III-A Partially Observable Markov Decision Process
We model adversarial search and tracking as a Partially Observable Markov Decision Process (POMDP). A POMDP for an agent can be defined by a state $s$, private observation $o$, action $a$, a transition function $\mathcal{T}:s\times a\mapsto s$. At each timestep $t$, an agent receives an observation $o^{t}$, chooses an action $a^{t}\in\mathcal{A}$, and obtains a reward $r^{t}:s\times a_{i}\mapsto\mathbb{R}$. The initial world state is drawn from a prior distribution $\rho$. In our environment (see §III-B), we assume the evader starts from a random position from in a corner region of the map while the pursuit team starts from a random position on the map.
III-B Prisoner Escape and Narco Interdiction Domains
In our evaluation of the proposed algorithm, we adjust two large-scale pursuit-evasion domains from prior work [14] named Prisoner Escape and Narco Interdiction (Figure 2). In both domains, an evader (prisoner or smuggler) is being tracked by a team of search agents in a partially observable environment. We draw three hideouts each episodes from all the candidate hideouts. The game-ending goal of the evader is to reach any hideout on the map (considered as success), or reach the time limit (considered as failure). The heterogeneous pursuit team consists of static (cameras) and dynamic agents (search parties and helicopters) that collaborate to search and track the evader in both domains.
The maximum speed of the evader is always lower than the pursuit agents, making it crucial for the evader to plan evasive maneuvers to successfully reach the hideouts undetected. The observation of the evader includes the current timestep, hideout locations, mountain locations, pursuer states and evader states. However, the camera locations and visibility map are hidden to the evader. These domains increase the complexity of the task and more closely match real-world dynamics compared to other pursuit-evasion games.
III-C Pursuit Team Behavior
We create a heuristic policy for tracking the evader inspired by previous object searching works [49, 50] and the fact that pursuit teams typically follow some routine patterns. The search agents can choose to either 1) go to the last known location of the evader, 2) intercept the evader, or 3) search in the vicinity of the evader’s last known location. We assume the detection range of searching team $d_{s}$ is parameterized by the agent type $\beta_{type}$, visibility $c_{v}$ and evader speed $s_{e}$: $d_{s}=\alpha\cdot(\beta_{type}\cdot c_{v}\cdot s_{e}+\eta)$, which means a higher visibility and evader speed will lead to higher detection range. The pursuer policy is designed to effectively utilize the available information and adapt to a wide variety of different scenarios.
IV Method
The evader’s objective is to reach one of the hideouts on the map and evade detection by the pursuit team, using only partial map knowledge and lacking any information about camera locations or models for whether the pursuers are able to detect the evader. RL or diffusion-based motion planning each partially fits our setting: RL can hypothetically learn to evade through trial-and-error. However, RL has difficulties with exploration in the large, adversarial, and multi-goal domains. On the other hand, diffusion-based motion planning can generate diverse plans that reach the goals; however, diffusion models struggle to learn evasive behaviors and avoid the pursuing agents. This weaknesses is due to the diffusion model is trained with a denoising loss objective that is not readily incorporated into RL [51]. Additionally, the partial observability induces high variance in training.
We propose to reap the benefits of both methods by leveraging diffusion models to guide RL exploration during training and generate path candidates at inference. We break down the hierarchy into the following: learning diffusion-based global plans that meet start, terminal, and obstacle constraints from RRT* paths (§IV-A), employing RL local evasive behaviors to avoid detection on route (§IV-B), and path selection from reversely constructed costmap during inference (§IV-C).
IV-A Diffusion Based Global Plan
In this section, we describe how we leverage diffusion models as a generative motion planner to provide a prior for the RL agent to traverse in large, partially observable, and adversarial environments [52]. Since A* cannot generate the optimal path to guide RL exploration on partially observable maps with adversarial agents[53], we utilize a diffusion model to plan waypoints for an approximate path that the agent should follow. By using the diffusion model, our method is sample-time efficient and enables the downstream reinforcement learning algorithm to outperform prior approaches. The underlying insight is to produce diverse options from the diffusion model for global exploration to deal with map uncertainty.
The diffusion path planning should have three key attributes: 1) global constraints are satisfied (start and terminal locations, obstacles etc.); 2) generated plans are diverse and multimodal and 3) the sampling time of the diffusion model is low. We guide the diffusion sampling procedure using constraint guidance [15] to best satisfy the constraints and use a small number of sampling timesteps to keep the sampling time low. This low sampling time allows us to run the diffusion model during the RL training procedure which is difficult for traditional planners (e.g., RRT* and A*). Additionally, the parallel sampling at inference time provides diverse candidates for the agent to use.
Data: Start Location $\mathcal{S}$, Goal Locations $\mathcal{G}$, Obstacle Location $\mathcal{M}$, Constraints $\mathcal{C}=<\mathcal{S},\mathcal{G},\mathcal{M}>$, Data buffer $\mathcal{D}$, Noise Scheduling Terms $\alpha_{i},\bar{\alpha_{i}}$, Learning Rate $\alpha$, Denoising Timesteps $T$
Result: Diffusion Model $s_{\theta}$
1 while buffer $\mathcal{D}$ not full do
       $c=<p_{s},p_{g},m>\sim\mathcal{C}$ ;
[…]
       $\tau\leftarrow RRT^{*}(c)$ ;
[…]
       $\tau\leftarrow Sample_{\downarrow}(\tau)$ ;
[…]
       $\mathcal{D}\leftarrow\tau,c$ ;
[…]
2
3while not converged do
[…]
4       $\tau_{0},c\sim\mathcal{D}$$i\sim\mathcal{U}(1,T),\epsilon\sim\mathcal{N}(0,I)$;
[…]
5       $\tau_{i}\leftarrow\sqrt{\bar{\alpha}_{i}}\tau_{0}+\sqrt{1-\bar{\alpha}}_{i}\epsilon$;
6       $\mathcal{L}(\theta)=\|\epsilon-s_{\theta}(\tau_{i},i,c))\|_{2}^{2}$;
[…]
7       $\theta=\theta+\alpha\nabla_{\theta}\mathcal{L}(\theta)$
Algorithm 1 Diffusion Model Training
We outline the diffusion training process in Algorithm 1. We first generate paths with RRT* (Line 1) — a sample-based motion planner that yields varied paths even for the same initial and target states, which enables the diffusion model to similarly generate diverse paths. Then we uniformly downsample the path into sparse waypoints (Line 1) and train a diffusion model based on these waypoints (Line 1-1). This process differs from Caravalho et al. [19], which trains a diffusion model directly from the dense and smoothed paths. Generating a path with a few waypoints is sufficient enough to act as the high level planner as we expect the RL policy to act as the low-level dynamic actor to move between waypoints. Generating fewer waypoints also allows sampling time to be faster, which is crucial for training in an RL loop.
Data: Constraints $\mathcal{C}$
Result: Global plan $\tau^{0}$
1 while $\tau^{0}$ not satisfies constraints $\mathcal{C}$ do
       $\tau^{T}\leftarrow$ sample from $\mathcal{N}(0,I)$ ;
[…]
2       for all $i$ from $T$ to $1$ do
3             $(\mu^{i},\Sigma^{i})\leftarrow s_{\theta}(\tau^{i}),\Sigma_{\theta}(\tau^{i})$
             $\tau^{i-1}\sim\mathcal{N}(\mu^{i},\Sigma^{i})$ ;
[…]
             $\tau^{i-1}\leftarrow\mathcal{C}(\tau^{i-1})$ ;
[…]
4
5
Return $\tau^{0}$
Algorithm 2 Diffusion Path Generator $\pi^{d}$
The sampling process is similar to our previous work [15] summarized in Algorithm 2 as diffusion path generator $\pi^{d}$. We apply global constraints at each denoising step (Line 2) to help generate a valid path from start to the end location.
IV-B RL Based Adversarial Behavior Learning
In addition to our diffusion model considering the global objectives, it is also important to design an algorithm directly interacting with the environment that can not only accomplish the tasks from the high-level planners but also capture the dynamic patterns in the map to achieve the local adversarial detection avoiding goals. Even the heuristics have been built in our previous works, we believe that a data-driven reinforcement learning algorithm is a more convenient way compared with the man-crafted policy and can perform better within a certain distribution.
Data: Initial state $s_{0}$, Diffusion Global Planner $\pi^{d}$, SAC replay buffer $\mathcal{D}_{s}$
Result: SAC low-level policy $\pi^{l*}$
1 while training episodes do
       $\{w_{i}\}_{i=0}^{N_{w}}=\mathbf{w}\sim\pi^{d}(\mathbf{w}|s_{0})$ ;
[…]
2       while $w_{i}$ not reached do
             $O_{aug}\leftarrow[O|w_{i}]$ ;
[…]
             $a\sim\pi^{l}(a|O_{aug})$ ;
[…]
             $O_{aug}^{\prime},r\leftarrow\mathbf{EnvStep}(a)$ ;
[…]
             $\mathcal{D}_{s}\leftarrow\{O_{aug},a,O_{aug}^{\prime},r\}$ ;
[…]
3             if $w_{i}$ reached then
                   $i\leftarrow i+1$ ;
[…]
4
5
      $\pi^{l*}\leftarrow SAC(d_{s}\sim\mathcal{D}_{s})$ ;
[…]
6
Algorithm 3 Adversarial Behavior Learning
Therefore we employ an RL algorithm Soft Actor-Critic (SAC) as the low-level policy with the following aims: 1) basically follow the waypoints from diffusion path and 2) learn to adjust its velocities to evade from searching agents. We summarize the RL training process in Algorithm 3 where we augment the observation with the current desired waypoint from the high-level planner (Line 3, 3). The training objective can be formulated as Equation (1) where $\approx$ means the similar low-level policy can be derived from both sides.
$R=\max_{\pi}\mathbb{E}_{\pi}$ $\left[\sum_{t}\gamma^{t}\cdot r\right]$
$\approx\max_{\pi^{g}}\mathbb{E}_{\pi^{g}}$ $\left[\max_{\pi^{l}}\mathbb{E}_{\pi^{l}}\left[\sum_{\tau=0}^{t_{g}}\sum_{t=0}^{t_{l}}\gamma^{t}\cdot r\bigg{|}w^{\tau}\sim\pi^{g}(w|s^{\tau})\right]\right]$
$\approx\max_{\pi^{l}}\mathbb{E}_{\pi^{l}}$ $\left[\sum_{\tau=0}^{t_{g}}\sum_{t=0}^{t_{l}}\gamma^{t}\cdot r\bigg{|}\mathbf{w}\sim\pi^{d}(\mathbf{w}|s_{0}),s_{0}\sim\rho\right]$ (1)
$\rho$ is the initial environment distribution for the start and hideout locations ($s_{0}\sim\rho$ is omitted from the derivation), $\pi^{d}$ is the diffusion global planner, and $\mathbf{w}$ is the generated waypoints to be tracked, $t_{g}$ is the number of waypoints. This equation shows that we are able to decompose our adversarial evading learning problem into a sequential optimization problem, including a waypoint generator and a low-level waypoint tracker, both of which aim to maximize the discounted accumulative reward. However, the unconstrained random exploration of these policies will subject to high variance from the partial observability, therefore we confine our exploration in a more reasonable region described by the diffusion policy distribution in the previous section. We set reward terms to guide the evader to reach the waypoint and avoid detection in Equation (2) where $r_{g}$ is the reward given when reaching the waypoint, $r_{adv}$ is the detection penalty, and $r_{d}$ is the distance penalty to the next waypoint.
$r=r_{g}+r_{d}+r_{adv}$ (2)
TABLE I: Diffusion-RL Benchmarks (mean $\pm$ std)
Domain Prisoner Escape Narco Interdiction
Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$
Non-Learning Approaches
A-Star Heuristic 0.721$\pm$0.189 0.344$\pm$0.238 0.980$\pm$0.140 0.635$\pm$0.295 0.392$\pm$0.305 0.800$\pm$0.400
RRT-Star Heuristic 0.731$\pm$0.176 0.321$\pm$0.220 0.940$\pm$0.237 0.655$\pm$0.308 0.375$\pm$0.319 0.830$\pm$0.376
VO-Heuristic 0.738$\pm$0.113 0.323$\pm$0.143 0.980$\pm$0.140 0.635$\pm$0.099 0.426$\pm$0.116 1.000$\pm$0.000
Learning Approaches
DDPG 0.234$\pm$0.342 0.725$\pm$0.404 0.070$\pm$0.255 0.305$\pm$0.321 0.645$\pm$0.374 0.000$\pm$0.000
SAC 0.779$\pm$0.117 0.058$\pm$0.108 0.130$\pm$0.336 0.486$\pm$0.347 0.434$\pm$0.406 0.010$\pm$0.099
Diffusion Only 0.779$\pm$0.092 0.276$\pm$0.116 1.000$\pm$0.000 0.776$\pm$0.127 0.261$\pm$0.148 1.000$\pm$0.000
Our Approaches
Diffusion-RL [ours] 0.899$\pm$0.087 0.116$\pm$0.095 0.960$\pm$0.196 0.886$\pm$0.106 0.123$\pm$0.102 0.940$\pm$0.237
Diffusion-RL-Map [ours] 0.941$\pm$0.079 0.068$\pm$0.080 0.980$\pm$0.140 0.943$\pm$0.069 0.066$\pm$0.080 1.000$\pm$0.000
IV-C Cost Map Construction and Inference
Even though we assume that the cameras and the terrain map information is totally unobservable for the evasive agent, we may know that there are some relations between the detection and these parts on the map. For example, the agent is easier to capture around the camera and more difficult to capture in the low visibility region. Therefore, it is possible for us to spatially show detection risks on a posterior task-oriented cost map associated with the policies we derived as a reverse engineering problem.
Algorithm 4 describes our procedure. We run a certain number of episodes after the hierarchy is trained well (Line 4) and simultaneously record the evader location $\mathbf{X}$ and its distance to the closest searching agent $\mathbf{d}$ (Line 4). Then we select out the evader locations $x_{i}$ with which the associated distance, $d_{i}$, is below a risk threshold $\epsilon$ and add a Gaussian distribution whose mean is $x_{i}$ and standard deviation $\sigma$ (Line 4-4). Next, we perform normalization and output the map.
Importantly, end-users can preview the global paths before execution or revise the cost map with the newly added global information, which adds more flexibility and explainability compared with traditional fully MDP based RL algorithms (Line 4). Finally we can select the global path with lowest cost based on the costmap in the inference (Line 4-4).
Data: Initial state $s_{0}$, Diffusion Global Planner $\pi^{d}$, Learned local policy $\pi^{l*}$, Costmap $\mathbf{G}$, Evasive agent location $\mathbf{x}$
Result: Selected global plan $\mathbf{w}^{*}$
1 while costmap constructing episodes do
[…]
2       $\mathbf{x},\mathbf{d}=\mathbf{Env}(s_{0},\pi^{d},\pi^{l*})$
3       for all $x_{i},d_{i}$ from $\mathbf{x},\mathbf{d}$ do
[…]
4             if $d_{i}<\epsilon$ then
                   $\mathbf{G}\leftarrow\mathbf{G}+\mathcal{N}(x_{i},\sigma^{2})$ ;
[…]
5
6
[…]
7       $\mathbf{G}\leftarrow adjust(\mathbf{G}/\mathbf{G}_{max})$;
[…]
8       $\mathcal{T}\leftarrow\{\mathbf{w}_{i}:\mathbf{w}_{i}\sim\pi^{d}(\mathbf{w}|s_{0})\}_{i=0}^{N_{s}}$ ;
[…]
10
Algorithm 4 Costmap and Inference
V Results & Discussion
In this section, we will first benchmark our diffusion-RL framework against both heuristics and learning based baselines. Then we visualize our constructed costmap which can help explain the results and show more advantages of our method. Then we justify our use of the diffusion model as the path generator. Next we validate our hierarchy in the real robot testbed. Finally we include a result discussion section.
V-A Achieving Long-Term Planning and Short-Term Evasion
We propose to use the following metrics to quantitatively evaluate our method and compare it against the baselines: detection, goal-reaching rate, and score ($\uparrow$ means higher is better and $\downarrow$ means lower is better in Table I). Detection is the average number of steps when the evader is detected for each episode. Goal-reaching rate is the number of episodes in which the evader reach the goal in the time limit divided by the total number of episodes tested. The score metric is a weighted reward term that balances the aforementioned two main objectives in our domains: reaching the goal ($\pm$50) and staying undetected by the opposing blue agents (-1 for each detection). Each metric is normalized by the min-max value of each category and we test the algorithms in 100 episodes with different random seeds from the training.
We compare our diffusion+RL method with learning based and non-learning based baselines. A* and RRT* heuristics are from our previous works [14] where the evader follows the A* or RRT* path and show some pre-scripted adversarial behaviors (e.g. heading to dense forest, changing velocity etc.) when the search team is nearby. Velocity obstacle (VO) method is widely used in the dynamic obstacle avoidance tasks which samples a velocity from the collision free velocity space [54]. From Table I, we find these heuristics can reach the goal at a high rate, but can be detected by the search team much more than our methods which leads to a lower score. The main reason is that the evader will be exposed to the unobservable cameras, which may attract the dynamic search team and the man-crafted evasive heuristic is not optimal for escaping. In comparison our approach, which utilizes the diffusion model as a motion planning module for a reinforcement learning agent, outperforms all non-learning baselines in terms of score and enables the agent to remain mostly undetected by the searching team while reaching the hideout (Table I).
Even though the reinforcement learning (DDPG & SAC) approaches directly optimize the score as their reward function, our methods outperform them on average by 109% on the score metric across both domains. At the same time, we find the goal reaching rate of the DDPG and SAC only methods are very low in prisoner escape domain (7%, 13%) and even lower in narco interdiction domain (0%, 1%). Interestingly, the SAC can outperform the DDPG in each metric we have across the two domains, which meets our expectation that SAC has more exploration ability from its stochastic policy and entropy regularization.
The SAC evader can be detected less than our map-based diffusion-RL algorithm by 19.12%, which is interesting and a red herring. Figure 3 shows a red colorbar for the distance between the evader and the closest agent in the pursuit team, and the diamond purple line indicates the diffusion planned global path. From the trajectory, we can see the SAC algorithm will lead the evader to the edge of the map where there is no camera but not the goal hideout. Instead, our method can loosely follow the diffusion waypoints (purple diamond) and simultaneously deviate a little to avoid the search team when they are close (indicated by magnet and orange dashed box). In addition, our method learns to reduce the evader speed when evading, which lowers the probability of detection, but maintains a higher speed when safer.
Figure 3: Our Diffusion-RL trajectory vs SAC trajectory. Colorbar indicates distance between evader and closest pursuer and purple line indicates the diffusion global path. We find that the our method (left) is able to evade the pursuers by deviating and decreasing the agent speed to reduce detection chance while still being able to reach the goal. On the contrary, the SAC policy is unable to reach the goal and only randomly changes its velocity.
V-B Costmap
(a) Costmap
(b) Camera Locations
(c) Selected Path
(d) Ad Hoc Map
(e) New Selected Path
Figure 4: Path Planning with the Costmap: The costmap (a) is constructed by correlating the agent’s risk of detection to its location on the map. We show in (b) that the agent is able to succesfully identify where the cameras are. Given this costmap, the agent can select a path that best evades high cost regions (c). Additional obstacles can be added ad-hoc (d) and a new path can be chosen (e). The grey areas indicate untraversable obstacles or danger zones.
As discussed in §IV-C, we build a costmap based on the distance between evader and pursuer which implies detection risks. The costmap serves as an interpretable and memory efficient way to allow our RL agent to plan based on past experiences. Figure 4a shows the costmap, where the high intensity region indicates the evader is close to the search team which may lead to a detection. We also find that the potential detection usually happens around the cameras (see Figure 4b), which is reasonable since the searching team will search the region after the evader is detected by the camera. Therefore, Figure 4c shows the evasive agent can select the best global path which receives lowest cost that can best conceal it from the camera array. At the inference stage, we can adjust the costmap with the new information we received. For example, we can add an undesired danger zone for the evader on the costmap (see Figure 4d) such that it will select another way to reach another hideout which will avoid the danger zone but potentially incurring more detection (see Figure 4e). We ablate our model with and without the costmap and show that the detection from map based path selection is further decreased by a half.
V-C Diffusion Paths Validation
The learned diffusion model should be able to generate diverse global paths that satisfy start and terminal constraints such that it can guide the low-level RL to explore high-valued state spaces. Additionally, the diffusion model should be sample-time efficient to accelerate RL training and inference. Figure 5a-5b compares the diffusion global paths trained on the datasets from RRT* and another commonly used path planner A*. We see the diffusion model encodes a more diverse distribution of the paths leading to the final hideouts when using RRT*. Figure 5c and Table 5d shows the planning time for both the diffusion and RRT* algorithms. The diffusion model takes 85.7% less time generating one trajectory compared with RRT*, which requires a map search. Furthermore, the difference is enlarged when planning more paths since we can draw samples from diffusion model in parallel. We generate the results using a machine with 11th Gen Intel Core i9 processor and GeForce GTX 1660 Ti graphic card.
(a) Diffusion Trajectories (A*)
(b) Diffusion Trajectories (RRT*)
(c) Time Comparison
Traj. RRT* Diffusion 1 0.7$\pm$0.2 0.1$\pm$0.0 10 6.2$\pm$0.3 0.2$\pm$0.0 20 12.5$\pm$0.6 0.4$\pm$0.1 30 19.7$\pm$1.5 0.4$\pm$0.0 40 26.3$\pm$1.9 0.5$\pm$0.0 50 33.0$\pm$2.3 0.6$\pm$0.2
(d) Time Comparison (mean±$\pm$±std)
Figure 5: The paths from the diffusion model trained on RRT* are more diverse than those trained on A* (5a-5b). Additionally, compared to the traditional RRT* planner, the diffusion model leverages the power of parallel computing to generate trajectories an order of magnitude faster (5c-5d).
V-D Real Robot Demonstration
We also test our method with the Robotarium [55] testbed to validate our method with real robot dynamics. Figure 6 shows the comparison between the RRT* heuristic and our method. We find the evader (red trajectory) equipped with our policy can reach the hideout and remain untracked by the searching team (blue and cyan trajectories) but the RRT* heuristic travels in a path with more cameras and is tracked by helicopter from $t\approx 90$. The result indicates that our data-driven learning based method can outperform the heuristics from the multi-agent interaction and map optimizations.
(a) RRT* Heuristic
(b) Diffusion-RL-Map (ours)
Figure 6: Pursuit-evasion trajectories: the evader, search party and helicopter trajectory are indicated with red, blue and cyan respectively. Evader is tracked from $t\approx 90$ with the RRT* heuristic but can escape from searching team with our method along the path to the hideout.
V-E Discussion
The significant disparity in goal-reaching rates between RL alone and our approaches indicates that RL struggles with long-horizon exploration in our domain, supporting our premise in §I that diffusion models can guide RL exploration. Additionally, Figure 3 reveals that SAC can become trapped in a local minimum. From this, we infer that our methods can achieve long-term planning by providing the RL agent with a diffusion-based global plan while retaining the agility to evade opponent agents in the short horizon. The cost map and time ablations in §V-B to §V-C demonstrate that our diffusion model can generate diverse samples and aid in downstream tasks such as cost map construction and path selection.
VI Conclusion and Future Work
In conclusion, we have presented a novel learning-based framework for evaders in large, partially observable, multi-agent pursuit-evasion settings. Our approach can help the evader reach the hideout and best conceal itself from capturing by using a hierarchy with a high-level diffusion path planner and a low-level RL evasive policy. Our work can also be extended to more general navigation problems (e.g. visual navigation, manipulator grasping etc.). However, our learned hierarchy requires additional interactions with the environments when constructing the costmap. In the future, we plan to integrate the map construction process into the RL training loop by iterating a location-only based value function $V(s)$ with the Bellman equation. Additionally, we will train the RL policy for both the pursuit and evasion sides simultaneously and investigate the relationship between the Nash equilibrium with respect to team composition.

### v2

@@
-Diffusion-Reinforcement Learning Hierarchical Motion Planning in Adversarial Multi-agent Games
+Diffusion-Reinforcement Learning Hierarchical Motion Planning in Multi-agent Adversarial Games
 Abstract
-Reinforcement Learning- (RL-)based motion planning has recently shown the potential to outperform traditional approaches from autonomous navigation to robot manipulation.
-In this work, we focus on a motion planning task for an evasive target in a partially observable multi-agent adversarial pursuit-evasion games (PEG).
-These pursuit-evasion problems are relevant to various applications, such as search and rescue operations and surveillance robots, where robots must effectively plan their actions to gather intelligence or accomplish mission tasks while avoiding detection or capture themselves.
-We propose a hierarchical architecture that integrates a high-level diffusion model to plan global paths responsive to environment data while a low-level RL algorithm reasons about evasive versus global path-following behavior.
-Our approach outperforms baselines by 51.2% by leveraging the diffusion model to guide the RL algorithm for more efficient exploration and improves the explanability and predictability.
-[^1] Code and implementation details are at https://github.com/CORE-Robotics-Lab/Opponent-Modeling
+Reinforcement Learning (RL)-based motion planning has recently shown the potential to outperform traditional approaches from autonomous navigation to robot manipulation.
+In this work, we focus on a motion planning task for an evasive target in a partially observable multi-agent adversarial pursuit-evasion game (PEG).
+Pursuit-evasion problems are relevant to various applications, such as search and rescue operations and surveillance robots, where robots must effectively plan their actions to gather intelligence or accomplish mission tasks while avoiding detection or capture.
+We propose a hierarchical architecture that integrates a high-level diffusion model to plan global paths responsive to environment data, while a low-level RL policy reasons about evasive versus global path-following behavior.
+The benchmark results across different domains and different observability show that our approach outperforms baselines by 77.18% and 47.38% on detection and goal reaching rate, which leads to 51.4% increasing of the performance score on average.
+Additionally, our method improves interpretability, flexibility and efficiency of the learned policy.
+[^1] Code and supplementary details are at https://github.com/ChampagneAndfragrance/Diffusion_RL
 I Introduction
-Multi-agent pursuit-evasion games (PEG) [1] are ubiquitous in real-world robotics scenarios where robotic systems could be applied to tracking drug smugglers [2], defensive escorting [3], or robot soccer [4].
+Multi-agent pursuit-evasion games (PEG) [1] are ubiquitous in real-world robotics scenarios, where robotic systems could be applied to drug smuggler tracking [2], defensive escorting [3], or robot soccer [4].
 In PEG, one team seeks to track its target while the target agent or team employs evasive maneuvers, and research naturally focuses on developing methods for either the pursuers [5] or evaders [6, 7, 8].
-While there has been extensive work on developing search policies, including robust tracking frameworks [9], topology map maintenance [10], and multi-hypothesis filters [11], relatively little attention has been paid to evasion.
-This lack of emphasis provides an opportunity to design a novel evasive policy which is also useful in the real-world scenarios (e.g.
-business ship evading from pirates).
-Traditional approaches to develop evasion policies typically assume full environment observability and are designed to optimize capture time [12] or an ad hoc utility function [13].
+Although there has been extensive work on developing search policies [9, 10, 11], relatively little attention has been paid to evasion.
+This lack of emphasis provides an opportunity to design a novel evasive policy that is useful in real-world scenarios, such as business ships evading pirates and evacuating lost people in national parks.
 Figure 1: Diffusion-RL Framework Overview: We first collect RRT* paths into a dataset.
@@
 A posterior costmap is built based on the learned hierarchy and detection risk and can be used to select the best global path in the inference stage.
-More recently, learning-based methods, such as Graph Neural Networks [14], diffusion models [15] or reinforcement learning (RL) methods [16] have been widely used in the pursuit-evasion games and in similar navigation problems [17, 18, 19] (e.g.
+Recently, learning-based methods, such as Graph Neural Networks [12], diffusion models [13] or reinforcement learning (RL) methods [14] have been widely used in the pursuit-evasion games and similar navigation problems [15, 16, 17] (e.g.
 goal-reaching, path-planning, etc.).
 These methods often outperform expert policies [18].
-Unfortunately, learning the RL policy is usually subject to two key challenges—partial observability and high sample complexity when exploring in large-scale, partially observable environments, as considered in our work.
+However, learning the RL policy is usually subject to two key challenges—partial observability and high sample complexity when exploring large-scale, partially observable environments, as considered in our work.
 Researchers propose hierarchical approaches to help exploration by structuring the goal, policies, or rewards.
-Goal-oriented reinforcement learning (e.g.
-hindsight experience replay [21]) increases the reward density by assuming the goal has been reached in each exploration trajectory and correcting observation and reward accordingly.
-Hierarchical reinforcement learning [22] improves the sample efficiency from spatial or temporal abstraction by multiple hierarchically organized RL policies, and hybrid reward structure based methods [23] try to capture the multi-modal distribution of reward where each part of value function can be a low-dimensional representation.
-However, their applications are either tested in fully observable games [22] or single agent scenarios [24], and sometimes require expert prior knowledge as primitives [25].
-In this paper, we propose a novel hierarchical learning framework that utilizes diffusion and RL as high and low level policies respectively.
+Goal-oriented reinforcement learning (for example, a replay of the hindsight experience [19]) increases the reward density by assuming that the goal has been reached in each exploration trajectory and corrects observation and reward accordingly.
+Hierarchical RL [20] improves sample efficiency from spatial or temporal abstraction by multiple hierarchically organized RL policies, and hybrid reward structure-based methods [21] try to capture the multimodal distribution of reward where each part of the value function can be a low-dimensional representation.
+However, their applications are tested in fully observable games [20] or require expert prior knowledge as primitives [22].
+In this paper, we propose a novel hierarchical learning framework that combines diffusion and RL as high- and low-level policies, respectively.
 The diffusion model is used as a global path generator to satisfy static map constraints, such as terminal states and obstacle avoidance.
-Additionally, we employ a low-level RL policy that learns to follow the waypoints while prudently evading captures (see Figure 1).
-Our key insight is that using the diffusion model can improve the RL performance by constraining its exploration to high-value state-action regions.
+Furthermore, we employ a low-level RL policy that learns to follow the waypoints while prudently evading captures (see Figure 1).
+Our key insight is that using the diffusion model can improve the RL performance by constraining its exploration to high-value state-action regions and can help downstream tasks by providing diverse options.
 We find our method is especially useful for the evader to learn escaping behaviors in the large multi-agent, multi-goal partially observable settings — Prisoner Escape and Narco Interdiction Domain (see §III-B).
-Increasing understanding of potential evader strategies will help future development of countermeasures.
 Contributions: In summary, our contributions are three-fold:
   1. 1.
-    Our work is the first to create a learning-based framework for evaders in large partially observable environments that learns without expert domain knowledge.
+    We propose a novel hierarchical system consisting of a diffusion model as a high-level global path planner to aid RL exploration and a low-level RL agent to learn evasive maneuvers, which outperforms all baselines.
   2. 2.
-    We propose a novel hierarchical system consisting of a diffusion model as a high level global path planner and a low level RL agent to learn evasive maneuvers, which significantly outperforms all baselines.
+    We design an algorithm to implicitly infer the cost map that is used to select paths within our hierarchical motion planning framework to evade the searching agents.
+Multiple ablation studies show the interpretability, flexibility, efficiency, and generalizability of our method.
   3. 3.
-    To the best of our knowledge, we pioneer the use of task-oriented costmap construction to enhance explainability, predictability, and flexibility, thereby facilitating a deeper understanding of the evader’s performance.
+    Our work is the first to equip evaders with hierarchical motion planning to learn escaping while reaching the navigation goal in large partially observable environments without expert primitives.
 II Related Work
 II-A Pursuit-Evasion Games
-A perfect information PEG (i.e., full observability) can be modeled as a differential game and solved using control theory [26, 27, 28, 29].
-In comparison, the incomplete information PEG (i.e., partial observability) is a more realistic setting.
-Researchers commonly use probabilistic models [30, 31] and formulate PEG as optimization problems with the objective of maximizing the payoff [32, 33].
-Therefore, recent research also focuses on deriving the optimal policy via RL [34, 32, 35].
-Unfortunately, most of RL work focuses on pursuit policies [34, 32, 36].
-While Generative Adversarial Imitation Learning has been employed for training an evader policy, this approach requires a dataset showing expert evasive behaviors and a complex curriculum RL scheme [35].
-Furthermore, prior work only validates this approach on a small map.
-In comparison, our method trains evasive RL policy from scratch without any evasion behavior demonstrations and is simpler and more suitable for large, adversarial, and partially observable domains.
+A perfect information PEG (i.e., full observability) can be modeled as a differential game and solved using control theory [23].
+In comparison, the incomplete information PEG (i.e., partial observability) is a more realistic setting since agents usually do not have full knowledge of the environment at all time.
+Researchers commonly use probabilistic models [24, 25] and formulate PEG as optimization problems to maximize the payoff [26, 27].
+Therefore, recent research also focuses on obtaining the optimal policy through RL [28].
+Unfortunately, most RL works focus on pursuit policies [29, 30].
+Some recent works [31, 32] have focused on evasive behaviors evolved from scratch, but these policies do not need to reach explicit navigation goals.
+In comparison, we derive an evader policy that can not only escape from pursuers but also reach goals in large, adversarial, and partially observable domains.
 II-B Diffusion Models in Path Generating
-Diffusion models are a class of generative models that have demonstrated impressive performance in generating sequential data across several applications, such as image synthesis and video generation [37, 38, 39].
+Diffusion models are a class of generative models that have demonstrated impressive performance in generating sequential data in several applications, such as image synthesis and video generation [33, 34, 35].
 Recently, there have been several works employing diffusion models in trajectory generation tasks.
 Diffuser [36] trains a diffusion model for trajectory generation on offline datasets and plans future trajectories with guided sampling.
-Diffusion policy extended this work to imitation learning and was able to learn robust policies for various pushing tasks conditioned only on visual inputs [41].
+The diffusion policy extends this work to imitation learning and demonstrates the ability to learn robust policies for various pushing tasks conditioned only on visual input [37].
 Recent work [13] has extended these approaches for generating trajectories of multiple opponents.
-There have also been many following works using diffusion models to synthesize trajectories and augment experience buffer for efficient offline RL training [42, 43].
-However, these works only function in fully-observable domains, where building a world model is feasible.
-We instead use the diffusion model as a trajectory planner that allows the reinforcement learning agent to generate diverse plan options in our partially observable pursuit-evasion domain.
-II-C Reinforcement Learning in Motion Planning
-Recent works have demonstrated the success of RL for motion planning problems in dynamic environments [44, 45, 46, 47, 17].
-Many papers discuss how to use RL to avoid obstacles by analyzing exteroceptive sensor information such as RGB cameras [44, 47], LIDARs [46, 47], or laser data converted from images [45].
-However, none of these prior works assumes the an adversarial environment, where the “obstacles” are intelligent pursuers, and the map size is as large as in our scenario.
-The most similar work from the system design perspective is from [17] which also employs a hierarchical structure in goal-conditioned planning.
-However, their environment does not incorporate adversaries or multiple goals.
-Additionally, they require expert datasets for training their RL policy.
-Dynamic obstacles are considered in [48]; however, these obstacles can only run on a pre-defined segment with a constant velocity, and all the environment objects are observable such that A* search can be employed effectively.
-In our scenario, we assume the cameras are unobservable to the evader, while the pursuit team can leverage these cameras to detect and track the evader.
-Therefore, we use a diffusion model to generate diverse global candidate paths and make it possible for RL to do useful exploration about the map.
+In our work, we further use the diffusion model as a trajectory planner that allows the reinforcement learning agent to generate diverse plan options in our partially observable pursuit-evasion domain.
+II-C RL in Motion Planning
+Recent work has demonstrated the success of RL for motion planning problems in dynamic environments [38, 39, 40, 15].
+Many papers discuss how to use RL to avoid obstacles by analyzing exteroceptive sensor information such as RGB cameras [40], LIDARs [39, 40], or converted laser data from images [38].
+However, none of these previous works assumes an adversarial environment where the “obstacles” are intelligent pursuers, or the map is as large as in our scenario with undetectable parts.
+The work [15] also employs a hierarchical structure in goal-conditioned planning; however, their environment does not incorporate adversaries or multiple goals, and their RL policy is not trained from scratch.
+Dynamic obstacles are considered in [41]; however, these obstacles can only run on a predefined segment with a constant velocity, and the map is known such that A* search can be sufficient to find the optimal path.
+In our scenario, we assume the cameras are hidden from the evader in a non-discovered terrain map, while the pursuit team can use these cameras to detect and track the evader.
 III Environment
@@
 We model adversarial search and tracking as a Partially Observable Markov Decision Process (POMDP).
-A POMDP for an agent can be defined by a state $s$, private observation $o$, action $a$, a transition function $\mathcal{T}:s\times a\mapsto s$.
+A POMDP for an agent can be defined by a state $s$, private observation $o$, action $a$, and a transition function $\mathcal{T}:s\times a\mapsto s$.
 At each timestep $t$, an agent receives an observation $o^{t}$, chooses an action $a^{t}\in\mathcal{A}$, and obtains a reward $r^{t}:s\times a_{i}\mapsto\mathbb{R}$.
 The initial world state is drawn from a prior distribution $\rho$.
-In our environment (see §III-B), we assume the evader starts from a random position from in a corner region of the map while the pursuit team starts from a random position on the map.
-III-B Prisoner Escape and Narco Interdiction Domains
-In our evaluation of the proposed algorithm, we adjust two large-scale pursuit-evasion domains from prior work [14] named Prisoner Escape and Narco Interdiction (Figure 2).
+In our environment (see §III-B), we assume that the evader starts from a random position from a corner region of the map while the pursuit team starts from random positions on the map.
+It is a partially observable environment for the evader in that 1) cameras are invisible to the evader, 2) the terrain map is not provided a prior, and 3) the evader may only have a limited detection range.
+III-B Prisoner Escape and Narco Interdiction
+In our evaluation of the proposed algorithm, we employ two open-sourced large-scale pursuit-evasion domains [12] named Prisoner Escape and Narco Interdiction (Figure 2).
 In both domains, an evader (prisoner or smuggler) is being tracked by a team of search agents in a partially observable environment.
-We draw three hideouts each episodes from all the candidate hideouts.
-The game-ending goal of the evader is to reach any hideout on the map (considered as success), or reach the time limit (considered as failure).
-The heterogeneous pursuit team consists of static (cameras) and dynamic agents (search parties and helicopters) that collaborate to search and track the evader in both domains.
-The maximum speed of the evader is always lower than the pursuit agents, making it crucial for the evader to plan evasive maneuvers to successfully reach the hideouts undetected.
-The observation of the evader includes the current timestep, hideout locations, mountain locations, pursuer states and evader states.
-However, the camera locations and visibility map are hidden to the evader.
-These domains increase the complexity of the task and more closely match real-world dynamics compared to other pursuit-evasion games.
+In each episode, three hideouts are drawn from all candidate hideouts.
+These hideouts are only visible to the evader and unknown to the pursuit team, as adversarial teams do not communicate intent.
+The terminal objective of the evader is to reach any hideout on the map (considered a success) or reach the time limit (considered a failure).
+Simultaneously, the evader should learn to avoid pursuit team detection.
+The heterogeneous pursuit team consists of static (cameras) and dynamic agents (search parties, helicopters, search boats, and planes) that collaborate to search and track the evader in both domains.
+The maximum speed of the evader is always lower than the pursuit agents, making it more challenging for the evader to plan evasive maneuvers to successfully reach the hideouts undetected.
+The observation of the evader includes the current timestep, hideout locations, mountain locations, pursuer states, and evader states.
+The evader can see the pursuer anywhere in the global-view mode and within a field-of-view in the local-view mode, and the camera locations and visibility map are always hidden from the evader.
+In the local-view mode, evader’s field-of-view, $d_{e}$, is parameterized by the searching agent type, $\beta_{type}$, visibility, $c_{v}$, and searching agent speed, $s_{s}$: $d_{e}=\alpha\cdot(\beta_{type}\cdot c_{v}\cdot s_{s}+\eta)$.
+These domains increase the complexity of the task and more closely match real-world dynamics compared to other pursuit-evasion games [42, 43, 23].
 III-C Pursuit Team Behavior
-We create a heuristic policy for tracking the evader inspired by previous object searching works [49, 50] and the fact that pursuit teams typically follow some routine patterns.
-The search agents can choose to either 1) go to the last known location of the evader, 2) intercept the evader, or 3) search in the vicinity of the evader’s last known location.
-We assume the detection range of searching team $d_{s}$ is parameterized by the agent type $\beta_{type}$, visibility $c_{v}$ and evader speed $s_{e}$: $d_{s}=\alpha\cdot(\beta_{type}\cdot c_{v}\cdot s_{e}+\eta)$, which means a higher visibility and evader speed will lead to higher detection range.
-The pursuer policy is designed to effectively utilize the available information and adapt to a wide variety of different scenarios.
+To provide a fair comparison between our evading algorithm and the baselines, we use a fixed heuristic policy for the pursuit team inspired by prior works [44, 45] that achieved good performance in the prisoner escape domain [14].
+The search agents can choose to either 1) go to the last detection, 2) search along the predicted path, or 3) search in the vicinity of the last detection.
+Searching team detection range, $d_{s}$, is parameterized by the agent type, $\beta_{type}$, visibility, $c_{v}$, and evader speed $s_{e}$: $d_{s}=\alpha\cdot(\beta_{type}\cdot c_{v}\cdot s_{e}+\eta)$, which means a higher visibility and evader speed will lead to larger detection range.
+The pursuer policy is designed to effectively utilize the available information and adapt to different scenarios.
 IV Method
-The evader’s objective is to reach one of the hideouts on the map and evade detection by the pursuit team, using only partial map knowledge and lacking any information about camera locations or models for whether the pursuers are able to detect the evader.
+The evader’s objective is to reach one of the hideouts on the map and evade detection by the pursuit team, using only partial map knowledge, without any information about camera locations or models for whether the pursuers can detect the evader.
 RL or diffusion-based motion planning each partially fits our setting: RL can hypothetically learn to evade through trial-and-error.
@@
 On the other hand, diffusion-based motion planning can generate diverse plans that reach the goals; however, diffusion models struggle to learn evasive behaviors and avoid the pursuing agents.
-This weaknesses is due to the diffusion model is trained with a denoising loss objective that is not readily incorporated into RL [51].
-Additionally, the partial observability induces high variance in training.
-We propose to reap the benefits of both methods by leveraging diffusion models to guide RL exploration during training and generate path candidates at inference.
-We break down the hierarchy into the following: learning diffusion-based global plans that meet start, terminal, and obstacle constraints from RRT* paths (§IV-A), employing RL local evasive behaviors to avoid detection on route (§IV-B), and path selection from reversely constructed costmap during inference (§IV-C).
-IV-A Diffusion Based Global Plan
-In this section, we describe how we leverage diffusion models as a generative motion planner to provide a prior for the RL agent to traverse in large, partially observable, and adversarial environments [52].
-Since A* cannot generate the optimal path to guide RL exploration on partially observable maps with adversarial agents[53], we utilize a diffusion model to plan waypoints for an approximate path that the agent should follow.
-By using the diffusion model, our method is sample-time efficient and enables the downstream reinforcement learning algorithm to outperform prior approaches.
-The underlying insight is to produce diverse options from the diffusion model for global exploration to deal with map uncertainty.
-The diffusion path planning should have three key attributes: 1) global constraints are satisfied (start and terminal locations, obstacles etc.); 2) generated plans are diverse and multimodal and 3) the sampling time of the diffusion model is low.
-We guide the diffusion sampling procedure using constraint guidance [15] to best satisfy the constraints and use a small number of sampling timesteps to keep the sampling time low.
-This low sampling time allows us to run the diffusion model during the RL training procedure which is difficult for traditional planners (e.g., RRT* and A*).
-Additionally, the parallel sampling at inference time provides diverse candidates for the agent to use.
+The reason is that the diffusion model has a relatively long denoising process that is difficult to handle environment dynamics.
+Additionally, partial observability induces high variance in training.
+We propose to leverage the strengths of diffusion and RL policies by using diffusion models to guide RL exploration during training and generate path candidates at inference.
+We decompose the hierarchy into the following: learning diffusion-based global plans that meet start, terminal, and obstacle constraints from RRT* paths (§IV-A), employing RL local evasive behaviors to avoid detection en-route (§IV-B), and path selection from posteriorly constructed costmap during inference (§IV-C).
 Data: Start Location $\mathcal{S}$, Goal Locations $\mathcal{G}$, Obstacle Location $\mathcal{M}$, Constraints $\mathcal{C}=<\mathcal{S},\mathcal{G},\mathcal{M}>$, Data buffer $\mathcal{D}$, Noise Scheduling Terms $\alpha_{i},\bar{\alpha_{i}}$, Learning Rate $\alpha$, Denoising Timesteps $T$
@@
 1 while buffer $\mathcal{D}$ not full do
+2
       $c=<p_{s},p_{g},m>\sim\mathcal{C}$ ;
@@
 […]
-2
-3while not converged do
+3
+4while not converged do
 […]
@@
 Algorithm 1 Diffusion Model Training
-We outline the diffusion training process in Algorithm 1.
-We first generate paths with RRT* (Line 1) — a sample-based motion planner that yields varied paths even for the same initial and target states, which enables the diffusion model to similarly generate diverse paths.
-Then we uniformly downsample the path into sparse waypoints (Line 1) and train a diffusion model based on these waypoints (Line 1-1).
-This process differs from Caravalho et al.
-[19], which trains a diffusion model directly from the dense and smoothed paths.
-Generating a path with a few waypoints is sufficient enough to act as the high level planner as we expect the RL policy to act as the low-level dynamic actor to move between waypoints.
-Generating fewer waypoints also allows sampling time to be faster, which is crucial for training in an RL loop.
 Data: Constraints $\mathcal{C}$
@@
 Algorithm 2 Diffusion Path Generator $\pi^{d}$
-The sampling process is similar to our previous work [15] summarized in Algorithm 2 as diffusion path generator $\pi^{d}$.
-We apply global constraints at each denoising step (Line 2) to help generate a valid path from start to the end location.
-IV-B RL Based Adversarial Behavior Learning
-In addition to our diffusion model considering the global objectives, it is also important to design an algorithm directly interacting with the environment that can not only accomplish the tasks from the high-level planners but also capture the dynamic patterns in the map to achieve the local adversarial detection avoiding goals.
-Even the heuristics have been built in our previous works, we believe that a data-driven reinforcement learning algorithm is a more convenient way compared with the man-crafted policy and can perform better within a certain distribution.
+IV-A Diffusion Based Global Plan
+In this section, we describe how we leverage diffusion models as a generative motion planner to provide a prior for the RL agent to traverse in large, partially observable, and adversarial environments [46].
+We utilize a diffusion model to clone the distribution from RRT* planner and generate waypoints of diverse candidate paths that the agent should follow approximately.
+This diffusion path planner guides the exploration, increases the sample-time efficiency, and improves the flexibility of the downstream RL algorithm.
+The underlying insight is to constrain the RL exploration at a promising distribution implicitly encoded by the diffusion model considering the map constraints.
+The diffusion path planning should have three key attributes: 1) global constraints are satisfied (start and terminal locations, obstacles, etc.); 2) generated plans are diverse and multimodal and 3) the sampling time of the diffusion model is low.
+We guide the diffusion sampling procedure using constraint guidance [13] to best satisfy the constraints and use a small number of sampling timesteps to keep the sampling time low.
+The low sampling time allows us to run the diffusion model during the RL training procedure, which is difficult for traditional search-based planners.
+Additionally, the parallel sampling reduces the overhead at inference time and provides diverse candidates for the agent to use.
+We outline the diffusion training process in Algorithm 1.
+We first generate paths with RRT* (Line 1) — a sample-based motion planner that yields varied paths even for the same initial and target states, which enables the diffusion model to similarly generate diverse paths.
+Then we uniformly downsample the path into sparse waypoints (Line 1) and train a diffusion model based on these waypoints (Line 1-1).
+Generating a path with a few waypoints is sufficient enough to act as the high-level planner as we expect the RL policy to act as the low-level dynamic actor to move between waypoints.
+Generating fewer waypoints also allows sampling time to be faster, which is crucial for training in an RL loop.
+The sampling process is summarized in Algorithm 2 with diffusion path generator, $\pi^{d}$.
+We apply global constraints at each denoising step (Line 2) to help generate a valid path from start to goal.
 Data: Initial state $s_{0}$, Diffusion Global Planner $\pi^{d}$, SAC replay buffer $\mathcal{D}_{s}$
@@
 Algorithm 3 Adversarial Behavior Learning
-Therefore we employ an RL algorithm Soft Actor-Critic (SAC) as the low-level policy with the following aims: 1) basically follow the waypoints from diffusion path and 2) learn to adjust its velocities to evade from searching agents.
-We summarize the RL training process in Algorithm 3 where we augment the observation with the current desired waypoint from the high-level planner (Line 3, 3).
-The training objective can be formulated as Equation (1) where $\approx$ means the similar low-level policy can be derived from both sides.
+IV-B RL Based Adversarial Behavior Learning
+In addition to our diffusion model considering the global objectives, it is also essential to design an algorithm directly interacting with the environment that can not only accomplish the tasks from the high-level planners but also capture the dynamic patterns in the map to achieve the local adversarial detection avoiding goals.
+Although evading heuristics exist [12], we believe that a data-driven reinforcement learning algorithm is a more convenient way compared with the hand-crafted policy and can perform better within specific environment distribution.
+Therefore, we employ an RL algorithm Soft Actor-Critic (SAC) as the low-level policy with the following aims: 1) approximately follow the waypoints from the diffusion path and 2) adapt the evader velocities to escape from searching agents.
+We summarize RL training in Algorithm 3 where we augment the observation with the current desired waypoint from the high-level planner (Line 3, 3).
+The training objective can be formulated as Equation (1) where $\approx$ means a similar low-level policy can be derived from both sides.
 $R=\max_{\pi}\mathbb{E}_{\pi}$ $\left[\sum_{t}\gamma^{t}\cdot r\right]$
@@
 $\rho$ is the initial environment distribution for the start and hideout locations ($s_{0}\sim\rho$ is omitted from the derivation), $\pi^{d}$ is the diffusion global planner, and $\mathbf{w}$ is the generated waypoints to be tracked, $t_{g}$ is the number of waypoints.
-This equation shows that we are able to decompose our adversarial evading learning problem into a sequential optimization problem, including a waypoint generator and a low-level waypoint tracker, both of which aim to maximize the discounted accumulative reward.
-However, the unconstrained random exploration of these policies will subject to high variance from the partial observability, therefore we confine our exploration in a more reasonable region described by the diffusion policy distribution in the previous section.
-We set reward terms to guide the evader to reach the waypoint and avoid detection in Equation (2) where $r_{g}$ is the reward given when reaching the waypoint, $r_{adv}$ is the detection penalty, and $r_{d}$ is the distance penalty to the next waypoint.
+Equation (1) shows that we can decompose our adversarial evading learning problem into a sequential optimization problem, including a waypoint generator and a low-level waypoint tracker, both of which aim to maximize the discounted accumulative reward.
+However, the unconstrained random exploration of these policies will be subject to high variance from the partial observability; therefore, we confine our exploration in a more reasonable region described by the diffusion policy distribution in the previous section.
+We set reward terms to guide the evader to follow the waypoint and avoid detection in Equation (2) where $r_{g}$ is the reward obtained when reaching the waypoint, $r_{adv}$ is the detection penalty, and $r_{d}$ is the distance penalty to the next waypoint.
 $r=r_{g}+r_{d}+r_{adv}$ (2)
-TABLE I: Diffusion-RL Benchmarks (mean $\pm$ std)
-Domain Prisoner Escape Narco Interdiction
-Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$
+TABLE I: Diffusion-RL benchmarks (mean $\pm$ std), G: global-view mode, L: local-view mode
+Domain Prisoner Escape (G) Narco Interdiction (G) Prisoner Escape (L) Narco Interdiction (L)
+Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$
 Non-Learning Approaches
-A-Star Heuristic 0.721$\pm$0.189 0.344$\pm$0.238 0.980$\pm$0.140 0.635$\pm$0.295 0.392$\pm$0.305 0.800$\pm$0.400
-RRT-Star Heuristic 0.731$\pm$0.176 0.321$\pm$0.220 0.940$\pm$0.237 0.655$\pm$0.308 0.375$\pm$0.319 0.830$\pm$0.376
-VO-Heuristic 0.738$\pm$0.113 0.323$\pm$0.143 0.980$\pm$0.140 0.635$\pm$0.099 0.426$\pm$0.116 1.000$\pm$0.000
+A-Star Heuristic 0.72$\pm$0.19 0.34$\pm$0.24 0.98$\pm$0.14 0.64$\pm$0.30 0.39$\pm$0.31 0.80$\pm$0.40 0.80$\pm$0.18 0.25$\pm$0.15 1.00$\pm$0.00 0.53$\pm$0.29 0.51$\pm$0.31 0.76$\pm$0.43
+RRT-Star Heuristic 0.73$\pm$0.18 0.32$\pm$0.22 0.94$\pm$0.24 0.66$\pm$0.31 0.38$\pm$0.32 0.83$\pm$0.38 0.77$\pm$0.11 0.29$\pm$0.14 1.00$\pm$0.00 0.59$\pm$0.26 0.46$\pm$0.30 0.84$\pm$0.37
+VO-Heuristic 0.74$\pm$0.11 0.32$\pm$0.14 0.98$\pm$0.14 0.64$\pm$0.10 0.43$\pm$0.12 1.00$\pm$0.00 0.73$\pm$0.12 0.33$\pm$0.15 0.99$\pm$0.10 0.64$\pm$0.10 0.43$\pm$0.11 1.00$\pm$0.00
 Learning Approaches
-DDPG 0.234$\pm$0.342 0.725$\pm$0.404 0.070$\pm$0.255 0.305$\pm$0.321 0.645$\pm$0.374 0.000$\pm$0.000
-SAC 0.779$\pm$0.117 0.058$\pm$0.108 0.130$\pm$0.336 0.486$\pm$0.347 0.434$\pm$0.406 0.010$\pm$0.099
-Diffusion Only 0.779$\pm$0.092 0.276$\pm$0.116 1.000$\pm$0.000 0.776$\pm$0.127 0.261$\pm$0.148 1.000$\pm$0.000
+DDPG 0.23$\pm$0.34 0.73$\pm$0.40 0.07$\pm$0.26 0.31$\pm$0.32 0.65$\pm$0.37 0.00$\pm$0.00 0.23$\pm$0.34 0.73$\pm$0.40 0.07$\pm$0.26 0.26$\pm$0.36 0.70$\pm$0.42 0.00$\pm$0.00
+SAC 0.78$\pm$0.12 0.06$\pm$0.11 0.13$\pm$0.34 0.49$\pm$0.35 0.43$\pm$0.41 0.01$\pm$0.10 0.73$\pm$0.12 0.10$\pm$0.15 0.05$\pm$0.22 0.51$\pm$0.40 0.41$\pm$0.47 0.00$\pm$0.00
+Diffusion Only 0.78$\pm$0.09 0.28$\pm$0.12 1.00$\pm$0.00 0.78$\pm$0.13 0.26$\pm$0.15 1.00$\pm$0.00 0.79$\pm$0.10 0.26$\pm$0.13 1.00$\pm$0.00 0.78$\pm$0.13 0.26$\pm$0.15 1.00$\pm$0.00
 Our Approaches
-Diffusion-RL [ours] 0.899$\pm$0.087 0.116$\pm$0.095 0.960$\pm$0.196 0.886$\pm$0.106 0.123$\pm$0.102 0.940$\pm$0.237
-Diffusion-RL-Map [ours] 0.941$\pm$0.079 0.068$\pm$0.080 0.980$\pm$0.140 0.943$\pm$0.069 0.066$\pm$0.080 1.000$\pm$0.000
+Diffusion-RL 0.90$\pm$0.09 0.12$\pm$0.10 0.96$\pm$0.20 0.89$\pm$0.11 0.12$\pm$0.10 0.94$\pm$0.24 0.90$\pm$0.12 0.11$\pm$0.10 0.90$\pm$0.30 0.89$\pm$0.11 0.11$\pm$0.11 0.91$\pm$0.29
+Diffusion-RL-Map 0.94$\pm$0.08 0.07$\pm$0.08 0.98$\pm$0.14 0.94$\pm$0.07 0.07$\pm$0.08 1.00$\pm$0.00 0.95$\pm$0.09 0.05$\pm$0.07 0.95$\pm$0.22 0.94$\pm$0.08 0.06$\pm$0.07 0.95$\pm$0.22
 IV-C Cost Map Construction and Inference
-Even though we assume that the cameras and the terrain map information is totally unobservable for the evasive agent, we may know that there are some relations between the detection and these parts on the map.
-For example, the agent is easier to capture around the camera and more difficult to capture in the low visibility region.
-Therefore, it is possible for us to spatially show detection risks on a posterior task-oriented cost map associated with the policies we derived as a reverse engineering problem.
+Having discussed how the diffusion path planner could help the RL explore within the reward promising region in §IV-B and how the low-level SAC policy could be derived under the global path and initial environment distribution, we now take a step further and describe which global path sample should be drawn in each test case.
+Even though the evasive agent does not have direct access to the camera or the terrain map information, this agent can still perform reverse spatial inference on the map by interacting with the learned policies to identify regions of heightened risk.
+Following this idea, we construct a cost map to reflect the risk of detection which we can then use to score the global paths.
 Algorithm 4 describes our procedure.
 We run a certain number of episodes after the hierarchy is trained well (Line 4) and simultaneously record the evader location $\mathbf{X}$ and its distance to the closest searching agent $\mathbf{d}$ (Line 4).
-Then we select out the evader locations $x_{i}$ with which the associated distance, $d_{i}$, is below a risk threshold $\epsilon$ and add a Gaussian distribution whose mean is $x_{i}$ and standard deviation $\sigma$ (Line 4-4).
+If the searching agent is within the evader’s field of view, we select out the evader locations $x_{i}$ with which the associated distance, $d_{i}$, is below a risk threshold $\epsilon$ and add a Gaussian distribution whose mean is $x_{i}$ and standard deviation $\sigma$ (Line 4-4).
 Next, we perform normalization and output the map.
-Importantly, end-users can preview the global paths before execution or revise the cost map with the newly added global information, which adds more flexibility and explainability compared with traditional fully MDP based RL algorithms (Line 4).
-Finally we can select the global path with lowest cost based on the costmap in the inference (Line 4-4).
+Importantly, end-users can preview the global paths before execution or revise the cost map with the newly added global information, which adds more flexibility and interpretability compared with traditional fully MDP-based RL algorithms (Line 4).
+We can select the global path with the lowest cost based on the costmap (Line 4-4).
 Data: Initial state $s_{0}$, Diffusion Global Planner $\pi^{d}$, Learned local policy $\pi^{l*}$, Costmap $\mathbf{G}$, Evasive agent location $\mathbf{x}$
@@
 […]
-4             if $d_{i}<\epsilon$ then
+4             if $d_{i}<\epsilon$ see searching agent then
                    $\mathbf{G}\leftarrow\mathbf{G}+\mathcal{N}(x_{i},\sigma^{2})$ ;
@@
 V Results & Discussion
-In this section, we will first benchmark our diffusion-RL framework against both heuristics and learning based baselines.
-Then we visualize our constructed costmap which can help explain the results and show more advantages of our method.
-Then we justify our use of the diffusion model as the path generator.
-Next we validate our hierarchy in the real robot testbed.
-Finally we include a result discussion section.
-V-A Achieving Long-Term Planning and Short-Term Evasion
-We propose to use the following metrics to quantitatively evaluate our method and compare it against the baselines: detection, goal-reaching rate, and score ($\uparrow$ means higher is better and $\downarrow$ means lower is better in Table I).
-Detection is the average number of steps when the evader is detected for each episode.
-Goal-reaching rate is the number of episodes in which the evader reach the goal in the time limit divided by the total number of episodes tested.
-The score metric is a weighted reward term that balances the aforementioned two main objectives in our domains: reaching the goal ($\pm$50) and staying undetected by the opposing blue agents (-1 for each detection).
-Each metric is normalized by the min-max value of each category and we test the algorithms in 100 episodes with different random seeds from the training.
-We compare our diffusion+RL method with learning based and non-learning based baselines.
-A* and RRT* heuristics are from our previous works [14] where the evader follows the A* or RRT* path and show some pre-scripted adversarial behaviors (e.g.
-heading to dense forest, changing velocity etc.) when the search team is nearby.
-Velocity obstacle (VO) method is widely used in the dynamic obstacle avoidance tasks which samples a velocity from the collision free velocity space [54].
-From Table I, we find these heuristics can reach the goal at a high rate, but can be detected by the search team much more than our methods which leads to a lower score.
-The main reason is that the evader will be exposed to the unobservable cameras, which may attract the dynamic search team and the man-crafted evasive heuristic is not optimal for escaping.
-In comparison our approach, which utilizes the diffusion model as a motion planning module for a reinforcement learning agent, outperforms all non-learning baselines in terms of score and enables the agent to remain mostly undetected by the searching team while reaching the hideout (Table I).
-Even though the reinforcement learning (DDPG & SAC) approaches directly optimize the score as their reward function, our methods outperform them on average by 109% on the score metric across both domains.
-At the same time, we find the goal reaching rate of the DDPG and SAC only methods are very low in prisoner escape domain (7%, 13%) and even lower in narco interdiction domain (0%, 1%).
-Interestingly, the SAC can outperform the DDPG in each metric we have across the two domains, which meets our expectation that SAC has more exploration ability from its stochastic policy and entropy regularization.
-The SAC evader can be detected less than our map-based diffusion-RL algorithm by 19.12%, which is interesting and a red herring.
-Figure 3 shows a red colorbar for the distance between the evader and the closest agent in the pursuit team, and the diamond purple line indicates the diffusion planned global path.
-From the trajectory, we can see the SAC algorithm will lead the evader to the edge of the map where there is no camera but not the goal hideout.
-Instead, our method can loosely follow the diffusion waypoints (purple diamond) and simultaneously deviate a little to avoid the search team when they are close (indicated by magnet and orange dashed box).
-In addition, our method learns to reduce the evader speed when evading, which lowers the probability of detection, but maintains a higher speed when safer.
-Figure 3: Our Diffusion-RL trajectory vs SAC trajectory.
-Colorbar indicates distance between evader and closest pursuer and purple line indicates the diffusion global path.
-We find that the our method (left) is able to evade the pursuers by deviating and decreasing the agent speed to reduce detection chance while still being able to reach the goal.
-On the contrary, the SAC policy is unable to reach the goal and only randomly changes its velocity.
-V-B Costmap
+In this section, we will first benchmark our diffusion-RL framework for the global- and local-view evader against both heuristics and learning based baselines (see §V-A).
+Then we visualize our constructed costmap and show the interpretability, efficiency, and generalizability of our method in multiple ablation studies (see §V-BV-D).
+Next, we validate our hierarchy in the real robot testbed (see §V-E).
+Finally, we include a result discussion section (see §V-F).
+Figure 3: Our Diffusion-RL trajectory vs SAC trajectory. Colorbar indicates the distance between the evader and closest pursuer and the purple line indicates the diffusion global path.
 (a) Costmap
@@
 Figure 4: Path Planning with the Costmap: The costmap (a) is constructed by correlating the agent’s risk of detection to its location on the map.
-We show in (b) that the agent is able to succesfully identify where the cameras are.
-Given this costmap, the agent can select a path that best evades high cost regions (c).
+We show in (b) that the agent can successfully identify where the cameras are.
+Given this costmap, the agent can select a path that best evades high-cost regions (c).
 Additional obstacles can be added ad-hoc (d) and a new path can be chosen (e).
 The grey areas indicate untraversable obstacles or danger zones.
+V-A Achieving Long-Term Planning and Short-Term Evasion
+We propose to use the following metrics to quantitatively evaluate our method and compare it against the baselines: detection, goal-reaching rate, and score ($\uparrow$ means higher is better and $\downarrow$ means lower is better in Table I).
+Detection is the average number of steps when the evader is detected for each episode.
+Goal-reaching rate is the number of episodes in which the evader reaches the goal within the time limit divided by the total number of episodes tested.
+The score metric is a weighted reward term that balances the two main objectives in our domains: reaching the goal ($\pm$50) and staying undetected by the opposing blue agents (-1 for each detection).
+Each metric is normalized by the min-max value of each category and we test the algorithms in 100 episodes with different random seeds from the training.
+We compare our diffusion+RL method with learning based and non-learning-based baselines.
+A* and RRT* heuristics are from the open-sourced environment [12] where the evader follows the A* or RRT* path and shows some pre-scripted adversarial behaviors (e.g.
+heading to dense forest, changing velocity, etc.) when the search team is nearby.
+Velocity obstacle (VO) method is widely used in dynamic obstacle avoidance tasks which samples a velocity from the collision-free velocity space [47].
+From Table I, we find these heuristics can reach the goal at a high rate in two view modes across both domains.
+However, the heuristic evader is detected by the search team much more than our methods, which leads to a lower score.
+The main reason is that the evader will be exposed to the cameras hidden from it, which may attract the dynamic search team and the hand-crafted evasive heuristic is not optimal for escaping.
+Although RL-only approaches (DDPG [48] & SAC [49]) directly optimize the score as the reward function, our methods outperform them on average by 108% on the score metric across both domains.
+The SAC evader can be detected less than our diffusion-RL algorithm by 10.26% in prisoner escape domain, which is interesting and a red herring.
+Figure 3 shows a red colorbar for the distance between the evader and the closest agent in the pursuit team, and the diamond purple line indicates the diffusion planned global path.
+From the trajectory, we can see the SAC algorithm will lead the evader to the edge of the map where there is no camera but not the goal hideout.
+Instead, our method can loosely follow the diffusion waypoints (purple diamond) and simultaneously deviate a little to avoid the search team when they are close (indicated by magnet and orange dashed box).
+In addition, our method learns to reduce the evader speed when evading, which lowers the probability of detection, but maintains a higher speed when safer.
+We also ablate our diffusion-RL with the diffusion-only baseline (Diffuser [36]) and the 56.6% detection rate decreasing shows the low-level RL policy is crucial to adversarial evading behaviors.
+V-B Interpretability and Flexibility
 As discussed in §IV-C, we build a costmap based on the distance between evader and pursuer which implies detection risks.
-The costmap serves as an interpretable and memory efficient way to allow our RL agent to plan based on past experiences.
-Figure 4a shows the costmap, where the high intensity region indicates the evader is close to the search team which may lead to a detection.
+The costmap serves as an interpretable and memory-efficient way to allow our hierarchical agent to plan based on past experiences and will improve the flexibility in the inference stage compared with RL-only motion planning.
+Figure 4a shows the costmap, where the high-intensity region indicates the evader is close to the search team which may lead to detection.
 We also find that the potential detection usually happens around the cameras (see Figure 4b), which is reasonable since the searching team will search the region after the evader is detected by the camera.
-Therefore, Figure 4c shows the evasive agent can select the best global path which receives lowest cost that can best conceal it from the camera array.
+Therefore, Figure 4c shows the evasive agent can select the best global path which receives lowest cost and can best conceal it from the camera array.
 At the inference stage, we can adjust the costmap with the new information we received.
-For example, we can add an undesired danger zone for the evader on the costmap (see Figure 4d) such that it will select another way to reach another hideout which will avoid the danger zone but potentially incurring more detection (see Figure 4e).
-We ablate our model with and without the costmap and show that the detection from map based path selection is further decreased by a half.
-V-C Diffusion Paths Validation
-The learned diffusion model should be able to generate diverse global paths that satisfy start and terminal constraints such that it can guide the low-level RL to explore high-valued state spaces.
-Additionally, the diffusion model should be sample-time efficient to accelerate RL training and inference.
-Figure 5a-5b compares the diffusion global paths trained on the datasets from RRT* and another commonly used path planner A*.
-We see the diffusion model encodes a more diverse distribution of the paths leading to the final hideouts when using RRT*.
-Figure 5c and Table 5d shows the planning time for both the diffusion and RRT* algorithms.
-The diffusion model takes 85.7% less time generating one trajectory compared with RRT*, which requires a map search.
-Furthermore, the difference is enlarged when planning more paths since we can draw samples from diffusion model in parallel.
+For example, we can add an undesired danger zone for the evader on the costmap (see Figure 4d) such that it will select another way to reach another hideout which will avoid the danger zone (see Figure 4e).
+We ablate our model with and without the costmap and show that the detection from map-based path selection is further decreased by half.
+V-C Efficiency and Diversity in Guidance
+Figure 5a-5b compares the diffusion global paths trained on the datasets from RRT* and A*.
+We see the diffusion model encodes a more diverse distribution of the paths leading to the final hideouts.
+As a result, A* guided RL method can only receive a score of 0.78$\pm$0.09, which is worse than the performance of our diffusion-RL method (0.90$\pm$0.09).
+Table 5c shows the planning time in the inference stage for both the diffusion and RRT* algorithms.
+The diffusion model takes 85.7% less time to generate one trajectory compared with RRT*, which requires a map search.
+Furthermore, the difference is enlarged when planning more paths since we can draw samples from the diffusion model in parallel.
+In addition, Figure 5d shows that the diffusion-guided RL training uses less than a half time to reach the same reward level (save 25.2 hours), which greatly improves training time efficiency crucial to parameter tuning.
 We generate the results using a machine with 11th Gen Intel Core i9 processor and GeForce GTX 1660 Ti graphic card.
-(a) Diffusion Trajectories (A*)
-(b) Diffusion Trajectories (RRT*)
-(c) Time Comparison
+(a) A* Trajectories
+(b) Diffusion Trajectories
 Traj. RRT* Diffusion 1 0.7$\pm$0.2 0.1$\pm$0.0 10 6.2$\pm$0.3 0.2$\pm$0.0 20 12.5$\pm$0.6 0.4$\pm$0.1 30 19.7$\pm$1.5 0.4$\pm$0.0 40 26.3$\pm$1.9 0.5$\pm$0.0 50 33.0$\pm$2.3 0.6$\pm$0.2
-(d) Time Comparison (mean±$\pm$±std)
-Figure 5: The paths from the diffusion model trained on RRT* are more diverse than those trained on A* (5a-5b).
-Additionally, compared to the traditional RRT* planner, the diffusion model leverages the power of parallel computing to generate trajectories an order of magnitude faster (5c-5d).
-V-D Real Robot Demonstration
+(c) Inference Time (mean±$\pm$±std)
+(d) Training Time (Hours)
+Figure 5: The paths from the diffusion model trained on RRT* are more diverse than A* paths (5a-5b).
+Compared to the traditional RRT* planner, the diffusion model leverages the power of parallel computing to generate trajectories an order of magnitude faster (5c).
+The diffusion-guided RL training also significantly decreases the training time (5d).
+V-D Generailizability to New Camera Map
+The high-level diffusion and low-level RL policies trained on one map can be generalized to other maps with different camera distributions with the only need to re-run costmap reconstruction process rather than the whole training pipeline (see Figure 6).
+We benchmark with the new camera map, and our results are shown in the Table II.
+Our method still outperforms the best baseline by 18.6% and outperforms 43.99% across all baselines in average, which shows our algorithm can be generalized to the new camera map without any retraining of diffusion or RL policies.
+(a) New Camera Map
+(b) New Costmap
+Figure 6: (6a) - Cameras are highlighted.
+We use a different camera distribution in the test scenario (indicated by orange) from the one in the training scenario (indicated by blue).
+(6b) - we derive a new costmap reflecting the testing scenario camera distribution by running Algorithm 4.
+TABLE II: Generalizability Benchmarks (mean $\pm$ std)
+Domain (global-view) Prisoner Escape
+Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$
+Non-Learning Approaches
+A-Star Heuristic 0.694$\pm$0.185 0.376$\pm$0.229 0.970$\pm$0.171
+RRT-Star Heuristic 0.701$\pm$0.179 0.367$\pm$0.220 0.970$\pm$0.171
+VO-Heuristic 0.728$\pm$0.104 0.336$\pm$0.131 0.980$\pm$0.140
+Learning Approaches
+DDPG 0.136$\pm$0.241 0.847$\pm$0.264 0.070$\pm$0.255
+SAC 0.670$\pm$0.203 0.195$\pm$0.226 0.130$\pm$0.336
+Diffusion Only 0.767$\pm$0.080 0.292$\pm$0.100 1.000$\pm$0.000
+Our Approaches
+Diffusion-RL [ours] 0.864$\pm$0.111 0.155$\pm$0.116 0.940$\pm$0.237
+Diffusion-RL-Map [ours] 0.910$\pm$0.075 0.113$\pm$0.093 1.000$\pm$0.000
+V-E Real Robot Demonstration
 We also test our method with the Robotarium [50] testbed to validate our method with real robot dynamics.
-Figure 6 shows the comparison between the RRT* heuristic and our method.
-We find the evader (red trajectory) equipped with our policy can reach the hideout and remain untracked by the searching team (blue and cyan trajectories) but the RRT* heuristic travels in a path with more cameras and is tracked by helicopter from $t\approx 90$.
-The result indicates that our data-driven learning based method can outperform the heuristics from the multi-agent interaction and map optimizations.
+Figure 7 shows the comparison among the RRT* heuristic, SAC, and our method.
+We find the evader (red trajectory) equipped with our policy can reach the hideout and remain untracked by the searching team (blue and cyan trajectories).
+In comparison, the SAC policy does not lead the evader to the hideout, and the RRT* heuristic travels a path near cameras and is tracked by the helicopter starting at $t\approx 90$.
+This validates our benchmarking results and indicates that our hierarchical data-driven learning-based method can outperform the baselines from the efficient diffusion-guided exploration, RL-based trial-and-error multi-agent interaction, and map optimizations.
 (a) RRT* Heuristic
-(b) Diffusion-RL-Map (ours)
-Figure 6: Pursuit-evasion trajectories: the evader, search party and helicopter trajectory are indicated with red, blue and cyan respectively.
-Evader is tracked from $t\approx 90$ with the RRT* heuristic but can escape from searching team with our method along the path to the hideout.
-V-E Discussion
-The significant disparity in goal-reaching rates between RL alone and our approaches indicates that RL struggles with long-horizon exploration in our domain, supporting our premise in §I that diffusion models can guide RL exploration.
+(b) SAC
+(c) Diffusion-RL-Map
+Figure 7: Pursuit-evasion trajectories: the evader, search party, and helicopter trajectory are indicated with red, blue, and cyan, respectively.
+We use a diffusion model to guide the evader to the hideout and explore escaping behaviors with RL, which performs better than RRT* heuristic with insufficient evasive behaviors and SAC method converging to a suboptimal place.
+V-F Discussion
+The significant disparity in goal-reaching rates between standalone RL and our approaches indicates that RL struggles with long-horizon exploration in large partially observable domains, supporting our premise in §I that diffusion models can guide RL exploration.
 Additionally, Figure 3 reveals that SAC can become trapped in a local minimum.
 From this, we infer that our methods can achieve long-term planning by providing the RL agent with a diffusion-based global plan while retaining the agility to evade opponent agents in the short horizon.
-The cost map and time ablations in §V-B to §V-C demonstrate that our diffusion model can generate diverse samples and aid in downstream tasks such as cost map construction and path selection.
+The ablation studies in §V-B to §V-D demonstrate that our hierarchy can generate more interpretable paths, have more flexible motion plans and generalize to different test maps.
+Importantly, it significantly improves training and inference efficiency, which unlocks the potential on more applications in real-time tasks [51, 52].
 VI Conclusion and Future Work
@@
 Our approach can help the evader reach the hideout and best conceal itself from capturing by using a hierarchy with a high-level diffusion path planner and a low-level RL evasive policy.
-Our work can also be extended to more general navigation problems (e.g.
-visual navigation, manipulator grasping etc.).
-However, our learned hierarchy requires additional interactions with the environments when constructing the costmap.
-In the future, we plan to integrate the map construction process into the RL training loop by iterating a location-only based value function $V(s)$ with the Bellman equation.
-Additionally, we will train the RL policy for both the pursuit and evasion sides simultaneously and investigate the relationship between the Nash equilibrium with respect to team composition.
+In the future, we plan to train the whole framework end-to-end and bring our diffusion-RL technique to real-world navigation tasks involving high-level human domain knowledge transfer and low-level dynamic constraints (e.g.
+environment-aware path generation and finetuning in autonomous driving).